# Simultaneous Task Allocation and Planning for Multi-Robots under Hierarchical Temporal Logic Specifications

### v3

@@ VII Theoretic Analysis
Based on Def. VI.1, all automaton states within an initial search state are initial automaton states. By default, these are recognized as decomposition states. When a PBA ${\mathcal{P}}(r,\phi)$ is chosen for exploration, only the corresponding automaton state ${\mathbf{q}}(\phi)$ is updated, leaving the automaton states from other leaf specifications unaffected. The search can exit ${\mathcal{P}}(r,\phi)$ only through three types of switch transitions, each, by definition, associated with decomposition states. Upon switching to a different PBA ${\mathcal{P}}(r^{\prime},\phi^{\prime})$, the automaton state of $\phi$, which is a decomposition state, remains unchanged. Therefore, by inductive reasoning, it follows that at any given point, at most one automaton state from the leaf specifications does not belong to the decomposition set. ∎
@@
$|\to_{\mathcal{P}}|=\sum_{\phi\in\Phi^{K}}\sum_{r\in[N]}\left|\to_{{\mathcal{P}}(r,\phi)}\right|+\underbrace{|\zeta_{\text{in}}|+|\zeta_{\text{inter}}^{1}|+|\zeta_{\text{inter}}^{2}|}_{|\zeta|},$ (6)
which is the total number of in-spec transitions and switch transitions. According to [8], $|\to_{{\mathcal{P}}(r,\phi)}|$ is bounded by $|\to_{{\mathcal{P}}(r,\phi)}|\leq\left(|{\mathcal{Q}}_{\mathcal{B}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right)\left(|{\mathcal{Q}}_{\mathcal{B}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right).$ Furthermore, $|\zeta_{\text{in}}|$, $|\zeta_{\text{inter}}^{1}|$ and $|\zeta_{\text{inter}}^{2}|$ are, respectively, upper bounded by
$|\zeta_{\text{in}}|$ $\leq\sum_{\phi\in\Phi^{K}}\sum_{r\in[N-1]}\left(|{\mathcal{D}}_{\mathcal{B}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right)\left(|{\mathcal{D}}_{\mathcal{B}}(\phi)|\cdot|{\mathcal{S}}_{r+1}|\right),$
$|\zeta_{\text{inter}}^{1}|$ $\leq\sum_{\phi,\phi^{\prime}\in\Phi^{K}}\sum_{r\in[N]}\left(|{\mathcal{D}}_{\mathcal{B}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right)\left(|{\mathcal{D}}_{\mathcal{B}}(\phi^{\prime})|\cdot|{\mathcal{S}}_{r}|\right),$ (7)
$|\zeta_{\text{inter}}^{2}|$ $\leq\sum_{\phi,\phi^{\prime}\in\Phi^{K}}\sum_{r\in[N]}\left(|{\mathcal{Q}}^{F}_{\mathcal{B}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right)\left(|{\mathcal{D}}_{\mathcal{B}}(\phi^{\prime})|\cdot|{\mathcal{S}}_{1}|\right).$
@@
For a given PBA ${\mathcal{P}}(r,\phi)$, the number of states is upper bounded by $|{\mathcal{Q}}_{\mathcal{B}}(\phi)|\cdot|{\mathcal{S}}_{r}|$. Its squared term gives the upper bound for the number of in-spec transitions $|\to_{{\mathcal{P}}(r,\phi)}|$, since any pair of states within the PBA can potentially be connected. Regarding in-spec switch transitions between two consecutive PBAs, ${\mathcal{P}}(r,\phi)$ and ${\mathcal{P}}(r+1,\phi)$, each decomposition state can be associated with any robot state, which sets the upper bound on the number of possible starting states to $|{\mathcal{D}}_{\mathcal{B}}(\phi)|\times|{\mathcal{S}}_{r}|$. Similarly, the maximum number of end states is determined by $|{\mathcal{D}}_{\mathcal{B}}(\phi)|\times|{\mathcal{S}}_{r+1}|$, subsequently defining the upper bound for $|\zeta_{\text{in}}|$. Following the same logic, the upper bound for inter-spec switch transitions can be deduced as outlined in Eq. (VII.2). ∎
@@
The state-specification plan given by Alg. 2 satisfies hierarchical sc-LTL specifications $\Phi$.
@@
The soundness of Alg. 2 stems directly from its correct-by-construction design. To prove it, we use Alg. 1 to validate the correctness of the returned state-specification plan $\tau$. This plan $\tau$ is divided into segments $\tau=\tau_{1}\tau_{2}\ldots\tau_{l}$, with each $\tau_{k}$ aligning with a corresponding path segment $\Lambda_{k}$ from the searched path $\Lambda$. Beginning with the first segment $\tau_{1}$, and as the algorithm iterates over each $\phi\in\Phi^{K}\setminus\Phi^{acc}$ in line 1, the loop continues as long as $\phi\not=\phi_{1}$, where $\phi_{1}$ is the leaf specification associated with $\tau_{1}$. This is because $S(i)=\varnothing$ in line 1 for any $\phi\not=\phi_{1}$. When $\phi=\phi_{1}$, the set of reachable states ${\mathcal{Q}}_{\mathcal{B}}^{i+1}$ in NBA ${\mathcal{B}}(\phi_{1})$ are updated accordingly. Consequently, progress is only made for $\phi_{1}$, while all other leaf specifications remain idle. Upon completing the iteration over $\tau_{1}$ segment, $\phi_{1}$ reaches a decomposition state in NBA ${\mathcal{B}}(\phi_{1})$, as per the design of the hierarchical team models. In line 1 of Alg. 1, the non-leaf parent specifications of $\phi_{1}$ are updated based on the progress made solely in $\phi_{1}$. This update mirrors the operations performed by the UpdateNonLeafSpecs function in Alg. 4. The same procedure is applied to the subsequent segments $\tau_{2},\tau_{3},\ldots$, with each segment facilitating progress on a single leaf specification while pausing others. When an accepting state for a particular specification is reached, all PBAs associated with its child leaf specifications cease to be searched, as specified in Alg. 3, line 3. This is in alignment with the process in Alg. 1, line 1, where child leaf specifications of a satisfied parent specification are not further examined. Ultimately, upon reaching the end of $\tau_{l}$, the topmost specification is fulfilled, as noted in Alg. 2, line 2. This completion signifies that the hierarchical sc-LTL specifications are satisfied in accordance with Alg. 1, line 1, thereby confirming the correctness of the entire process. ∎
@@
Assuming that there exists a feasible state-specification plan $\tau=\tau^{0}\tau^{1}\ldots\tau^{h}$ that can be sequentially divided into $\ell$ segments with $\ell\leq|\Phi^{K}|$, represented as $\tau=\langle\tau\rangle_{1}\langle\tau\rangle_{2}\ldots\langle\tau\rangle_{\ell}$ where no two segments share the same specification. Each segment $\langle\tau\rangle_{l}$ fulfills a single leaf specification. That is, for every aggregated state-specification pair $\tau^{i}=((s_{1}^{i},\phi_{1}^{i}),(s_{2}^{i},\phi_{2}^{i}),\ldots,(s_{N}^{i},\phi_{N}^{i}))$ in the segment $\langle\tau\rangle_{l}=\tau^{m}\tau^{m+1}\ldots\tau^{n}$, each $\phi^{i}_{j}$ is either the same leaf specification $\phi$ that $\langle\tau\rangle_{l}$ satisfies, or it is the null specification $\epsilon$.
Note that the number of segments $\ell$ might be less than the number of leaf specifications $|\Phi^{K}|$, since it is possible that not all leaf specifications are required to be satisfied, such as leaf specifications that are connected by disjunction. The following properties only apply to Alg. 2 without heuristics.
@@
Assuming that a state-specification plan $\tau$ satisfying Asm. VII.5 exists, Alg. 2 is complete in that it is guaranteed to return a feasible plan that fulfills the hierarchical sc-LTL specifications $\Phi$.
@@
We proved the completeness of Alg. 2 by first disregarding inter-spec type I switch transitions. The inclusion of type I transitions merely enlarges the search space, without impacting the algorithm’s completeness. In this case, only inter-spec type II switch transitions exist between team models of any two leaf specifications $\phi$ and $\phi^{\prime}$. Type II transitions connect accepting states of every PBA within the product team model of $\phi$ (or $\phi^{\prime}$) to decomposition states of the first PBA of the team model of $\phi^{\prime}$ (or $\phi$). As a result, the search does not advance to product team models of other leaf specifications until it discovers a path segment that satisfies the current one, e.g., $\phi$. Furthermore, once an accepting state of a PBA for $\phi$ is reached, the search only moves to initial states, not other decomposition states, of the first PBA in another team model, e.g., $\phi^{\prime}$, as the search on this first PBA has not started yet.
@@
Assuming the existence of a state-specification plan $\tau^{*}$ that meets Asm. VII.5 with least cost, Alg. 2 is optimal in that the cost of the plan produced by Alg. 2 is equal to or less than the cost of $\tau^{*}$.
@@ VIII Simulation Experiments
We use Python 3.10.12 on a computer with 3.5 GHz Apple M2 Pro and 16G RAM. Our code can be available via the https://github.com/XushengLuo92/Hierarchical-LTL-STAP.
VIII-A User Study on Understanding Hierarchical sc-LTL
[…]
@@ VIII-A1 Survey
We carry out the survey on Prolific (www.prolific.com), recruiting 25 anonymous participants. The screening criteria specify that participants must have a STEM background and be at least undergraduate students, ensuring familiarity with mathematical logic. We begin the survey with an introduction to linear temporal logic, using the sorting task as a demonstration. Following this, participants answer two types of questions: subjective and objective. The subjective section consists of 7 questions where participants are presented with natural language instructions for 7 tasks. For each task, we provide a flat sc-LTL formula and hierarchical formulas. The participants are asked to choose which formula is easier to understand and clearer in representing the task described, of which task 5 is, “First place the circle, and then place diamond and star in any order, and lastly place the heart.” The flat sc-LTL is
@@
For the objective section, we prepare 4 pairs of flat and hierarchical formulas, each pair corresponding to the same task. This results in a total of 8 questions, with each one featuring either a flat sc-LTL formula or hierarchical sc-LTL formulas. We randomly order the formulas so that adjacent questions do not necessarily involve the same task. Participants are required to select from four natural language instructions the one that correspond to the given formula.
@@ VIII-A2 Results analysis
Responses with low accuracy on objective questions were excluded, indicating that respondents lack a solid understanding of temporal logics. This results in 20 valid surveys. For the subjective questions, we compute the percentage of participants who prefer flat versus hierarchical formulas across 140 instances (20 participants, 7 questions each). For the objective questions, we calculate the accuracy for each type of sc-LTL across 80 instances per sc-LTL type (20 participants, 4 questions per sc-LTL type). We track the time taken by each participant to respond to each question. A detailed analysis of the survey results is provided below.
For the subjective questions, 27.1% of all responses show a preference for flat sc-LTL, while 72.9% favor hierarchical sc-LTL. A detailed count of preferences for each task is illustrated in Fig. 6, where respondents showed a clear preference over hierarchical sc-LTL except for task 4. The specification is that “Place the circle, diamond, star and heart. Diamond should be after circle and heart should be after star.” The temporal constraints are simpler compared to those in task 5 in (8), making flat sc-LTL adequate for this scenario. Regarding the objective questions, the accuracy rates are 65.0% for flat sc-LTL and 86.3% for hierarchical sc-LTL. Additionally, the average response time for questions involving flat sc-LTL formulas is 78.3 seconds, compared to 47.1 seconds for hierarchical sc-LTL formulas. These findings further validate the user-friendliness of hierarchical sc-LTL, as evidenced by both higher accuracy and reduced response times. Details are presented in Fig. 7, where respondents achieved higher accuracy rates in all tasks specified under hierarchical sc-LTL.
@@ VIII-B Scenarios
We use the office environment described in [8] for service tasks, as depicted in Fig. 8. This 30×$\times 7$ grid map features 14 desk areas, 6 meeting rooms, and several other functional rooms. Additionally, we distribute $m$ robots at various locations within this simulated environment.
The first three scenarios are derived from the work [8]. We outline the hierarchical specifications for these scenarios below, while the corresponding flat versions of these specifications are referred to [8]. Note that the hierarchical specifications were derived directly from task descriptions, rather than from flat specifications. In Def. IV.1, we limit the sc-LTL specifications to co-safe LTL. However, all hierarchical sc-LTL specifications below contain a sub-formula expressed as $\square(a\Rightarrow b)$. Even with this sub-formula, it is still sufficient to assert satisfaction upon reaching the accepting states.
@@ VIII-B1 Scenario 1
Empty a paper bin located at desk $d_{5}$. The robot must avoid the public area while transporting the bin.
@@
$\phi_{2}^{2}=\Diamond(d_{5}\wedge\mathsf{emptybin}\wedge\bigcirc(d_{5}\wedge\mathsf{default}))$

### v4

@@ VII Theoretic Analysis
Based on Def. VI.1, all automaton states within an initial search state are initial automaton states. By default, these are recognized as decomposition states. When a PA ${\mathcal{P}}(r,\phi)$ is chosen for exploration, only the corresponding automaton state ${\mathbf{q}}(\phi)$ is updated, leaving the automaton states from other leaf specifications unaffected. The search can exit ${\mathcal{P}}(r,\phi)$ only through three types of switch transitions, each, by definition, associated with decomposition states. Upon switching to a different PA ${\mathcal{P}}(r^{\prime},\phi^{\prime})$, the automaton state of $\phi$, which is a decomposition state, remains unchanged. Therefore, by inductive reasoning, it follows that at any given point, at most one automaton state from the leaf specifications does not belong to the decomposition set. ∎
@@
$|\to_{\mathcal{P}}|=\sum_{\phi\in\Phi_{\text{leaf}}}\sum_{r\in[N]_{+}}\left|\to_{{\mathcal{P}}(r,\phi)}\right|+\underbrace{|\zeta_{\text{in}}|+|\zeta_{\text{inter}}^{1}|+|\zeta_{\text{inter}}^{2}|}_{|\zeta|},$ (7)
which is the total number of in-spec transitions and switch transitions. According to [9], $|\to_{{\mathcal{P}}(r,\phi)}|$ is bounded by $|\to_{{\mathcal{P}}(r,\phi)}|\leq\left(|{\mathcal{Q}}_{\mathcal{A}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right)\left(|{\mathcal{Q}}_{\mathcal{A}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right).$ Furthermore, $|\zeta_{\text{in}}|$, $|\zeta_{\text{inter}}^{1}|$ and $|\zeta_{\text{inter}}^{2}|$ are, respectively, upper bounded by
$|\zeta_{\text{in}}|$ $\leq\sum_{\phi\in\Phi_{\text{leaf}}}\sum_{r\in[N-1]_{+}}\left(|{\mathcal{D}}_{\mathcal{A}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right)\left(|{\mathcal{D}}_{\mathcal{A}}(\phi)|\cdot|{\mathcal{S}}_{r+1}|\right),$
$|\zeta_{\text{inter}}^{1}|$ $\leq\sum_{\phi,\phi^{\prime}\in\Phi_{\text{leaf}}}\sum_{r\in[N]_{+}}\left(|{\mathcal{D}}_{\mathcal{A}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right)\left(|{\mathcal{D}}_{\mathcal{A}}(\phi^{\prime})|\cdot|{\mathcal{S}}_{r}|\right),$ (8)
$|\zeta_{\text{inter}}^{2}|$ $\leq\sum_{\phi,\phi^{\prime}\in\Phi_{\text{leaf}}}\sum_{r\in[N]_{+}}\left(|{\mathcal{Q}}^{F}_{\mathcal{A}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right)\left(|{\mathcal{D}}_{\mathcal{A}}(\phi^{\prime})|\cdot|{\mathcal{S}}_{1}|\right).$
@@
For a given PA ${\mathcal{P}}(r,\phi)$, the number of states is upper bounded by $|{\mathcal{Q}}_{\mathcal{A}}(\phi)|\cdot|{\mathcal{S}}_{r}|$. Its squared term gives the upper bound for the number of in-spec transitions $|\to_{{\mathcal{P}}(r,\phi)}|$, since any pair of states within the PA can potentially be connected. Regarding in-spec switch transitions between two consecutive PAs, ${\mathcal{P}}(r,\phi)$ and ${\mathcal{P}}(r+1,\phi)$, each decomposition state can be associated with any robot state, which sets the upper bound on the number of possible starting states to $|{\mathcal{D}}_{\mathcal{A}}(\phi)|\times|{\mathcal{S}}_{r}|$. Similarly, the maximum number of end states is determined by $|{\mathcal{D}}_{\mathcal{A}}(\phi)|\times|{\mathcal{S}}_{r+1}|$, subsequently defining the upper bound for $|\zeta_{\text{in}}|$. Following the same logic, the upper bound for inter-spec switch transitions can be deduced as outlined in Eq. (VII.2). ∎
@@
The state-specification sequence given by Alg. 2 satisfies H-LTLf specifications $\Phi$.
@@
The soundness of Alg. 2 follows directly from its correct-by-construction design. The correctness of the state-specification sequence aligns with the correctness of the computed path $\Lambda=v_{0},v_{1},\ldots,v_{n}$. Each state $v\in\Lambda$ maintains its specification-related information through its component $v.{\mathbf{q}}$, which tracks the automaton states for all specifications. Consequently, for each specification $\phi\in\Phi$, we can derive an output word $\sigma(\phi)=\sigma_{0},\sigma_{1},\ldots,\sigma_{n}$, where $\sigma_{i}=\{\phi\}$ if $v_{i}.{\mathbf{q}}(\phi)\in{\mathcal{Q}}_{\mathcal{A}}^{F}(\phi)$ and $\sigma_{i}=\varnothing$ otherwise. Output words at lower levels propagate upward, generating those at the next higher level according to the bottom-up updates of non-leaf specifications in Alg. 4. Ultimately, the output word for the root specification $\phi_{1}^{1}$ includes $\phi_{1}^{1}$ itself, which meets the termination condition of Alg. 2 at line 2. As a result, the approach conforms to the semantic definition in Def. IV.6.
@@
Assuming that there exists a feasible state-specification sequence $\tau=\tau_{0}\tau_{1}\ldots\tau_{h}$ that can be sequentially divided into $\ell$ segments with $\ell\leq|\Phi_{\text{leaf}}|$, represented as $\tau=\langle\tau\rangle_{1}\langle\tau\rangle_{2}\ldots\langle\tau\rangle_{\ell}$ where no two segments share the same specification. Each segment $\langle\tau\rangle_{l}$ fulfills a single leaf specification. That is, for every aggregated state-specification pair $\tau_{i}=((s_{1}^{i},\psi_{1}^{i}),(s_{2}^{i},\psi_{2}^{i}),\ldots,(s_{N}^{i},\psi_{N}^{i}))$ in the segment $\langle\tau\rangle_{l}=\tau_{m}\tau_{m+1}\ldots\tau_{n}$, each $\psi^{i}_{r}$ is either the same leaf specification $\phi$ that $\langle\tau\rangle_{l}$ satisfies, or it is the null specification $\epsilon$.
Note that the number of segments $\ell$ might be less than the number of leaf specifications $|\Phi_{\text{leaf}}|$, since it is possible that not all leaf specifications are required to be satisfied, such as leaf specifications that are connected by disjunction. The following properties only apply to Alg. 2 without heuristics.
@@
Assuming that a state-specification sequence $\tau$ satisfying Asm. VII.5 exists, Alg. 2 is complete in that it is guaranteed to return a feasible plan that fulfills the H-LTLf specifications $\Phi$.
@@
We proved the completeness of Alg. 2 by first disregarding inter-spec type I switch transitions. The inclusion of type I transitions merely enlarges the search space, without impacting the algorithm’s completeness. In this case, only inter-spec type II switch transitions exist between team models of any two leaf specifications $\phi$ and $\phi^{\prime}$. Type II transitions connect accepting states of every PA within the product team model of $\phi$ (or $\phi^{\prime}$) to decomposition states of the first PA of the team model of $\phi^{\prime}$ (or $\phi$). As a result, the search does not advance to product team models of other leaf specifications until it discovers a path segment that satisfies the current one, e.g., $\phi$. Furthermore, once an accepting state of a PA for $\phi$ is reached, the search only moves to initial states, not other decomposition states, of the first PA in another team model, e.g., $\phi^{\prime}$, as the search on this first PA has not started yet.
@@
Assuming the existence of a state-specification sequence $\tau^{*}$ that meets Asm. VII.5 with least cost, Alg. 2 is optimal in that the cost of the plan produced by Alg. 2 is equal to or less than the cost of $\tau^{*}$.
@@ VIII Simulation Experiments
We use Python 3.10.12 on a computer with 3.5 GHz Apple M2 Pro and 16G RAM. Our code can be available via the https://github.com/XushengLuo92/Hierarchical-LTL-STAP. The video is accessible via the link https://www.youtube.com/watch?v=N3f8pUHDPF4&t=4s.
VIII-A User Study on Understanding H-LTLf
[…]
@@ VIII-A1 Survey
We carry out the survey on Prolific (www.prolific.com), recruiting 25 anonymous participants. The screening criteria specify that participants must have a STEM background and be at least undergraduate students, ensuring familiarity with mathematical logic. We begin the survey with an introduction to linear temporal logic, using the sorting task as a demonstration. Following this, participants answer two types of questions: subjective and objective. The subjective section consists of 7 questions where participants are presented with natural language instructions for 7 tasks. For each task, we provide a standard LTLf formula and hierarchical formulas. The participants are asked to choose which formula is easier to understand and clearer in representing the task described, of which task 5 is, “First place the circle, and then place diamond and star in any order, and lastly place the heart.” The standard LTLf is
@@
Note that standard LTLf and H-LTLf formulas are not equivalent, as achieving such equivalence is nontrivial. However, both formalisms are capable of capturing task descriptions to a large extent. For the objective section, we prepare 4 pairs of standard and hierarchical formulas, each pair corresponding to the same task. This results in a total of 8 questions, with each one featuring either a standard LTLf formula or H-LTLf formulas. We randomly order the formulas so that adjacent questions do not necessarily involve the same task. Participants are required to select from four natural language instructions the one that correspond to the given formula.
@@ VIII-A2 Results analysis
Responses with low accuracy on objective questions were excluded, indicating that respondents lack a solid understanding of temporal logics. This results in 20 valid surveys. For the subjective questions, we compute the percentage of participants who prefer standard versus hierarchical formulas across 140 instances (20 participants, 7 questions each). For the objective questions, we calculate the accuracy for each type of LTLf across 80 instances per LTLf type (20 participants, 4 questions per LTLf type). We track the time taken by each participant to respond to each question.
For the subjective questions, 27.1% of all responses show a preference for standard LTLf, while 72.9% favor H-LTLf. A detailed count of preferences is illustrated in Fig. 7, where respondents showed a clear preference over H-LTLf except for task 4. The specification is that “Place the circle, diamond, star and heart. Diamond should be after circle and heart should be after star.” The temporal constraints are simpler compared to those in task 5 in (9), making standard LTLf adequate for this scenario. Regarding the objective questions, the accuracy rates are 65.0% for standard LTLf and 86.3% for H-LTLf. Additionally, the average response time for questions involving standard LTLf formulas is 78.3 seconds, compared to 47.1 seconds for H-LTLf formulas. These findings further validate the user-friendliness of H-LTLf, as evidenced by both higher accuracy and reduced response times. Details are presented in Fig. 8, where respondents achieved higher accuracy rates in all tasks specified under H-LTLf.
@@ VIII-B Scenarios
We use the office environment described in [9] for service tasks, as depicted in Fig. 9. This 30×$\times 7$ grid map features 14 desk areas, 6 meeting rooms, and several other functional rooms. Additionally, we distribute $m$ robots at various locations. The first three scenarios are derived from the work [9]. We outline the hierarchical specifications for these scenarios below, while the corresponding standard versions of these specifications are referred to [9]. Note that the hierarchical specifications were derived directly from task descriptions, rather than from standard LTLf specifications.
@@ VIII-B1 Scenario 1
Transport the paper bin from desk $d_{5}$ to area $g$ for emptying, avoiding the public area while it is full. Return an empty bin from $g$ to desk $d_{5}$. The atomic propositions are: $\mathsf{default}$: the robot is not carrying any object, $\mathsf{carrybin}$: the robot is carrying a full paper bin, $\mathsf{dispose}$: the robot is disposing of garbage, $\mathsf{emptybin}$: the robot is carrying an empty paper bin, and $\mathsf{public}$: the robot is located in a public area.
@@
$\phi_{2}^{2}=\Diamond(g\wedge\bigcirc(g\wedge\mathsf{emptybin})\wedge\Diamond(d_{5}\wedge\bigcirc(d_{5}\wedge\mathsf{default})))$